OpenStudy (anonymous):

A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days: f(n) = 15(1.04)n Part A: When the scientist concluded his study, the height of the plant was approximately 20.53 cm. What is a reasonable domain to plot the growth function? (4 points) Part B: What does the y-intercept of the graph of the function f(n) represent? (2 points) Part C: What is the average rate of change of the function f(n) from n = 1 to n = 5, and what does it represent? I have no idea what I'm doing

myininaya (myininaya):

I think f(n)=15(1.04)^n right? the height is f(n) replace f(n) with 20.53 and solve for n 20.53=15(1.04)^n This will tell us the how many days (n) it will take for the height (f(n)) to be 20.53. Since this is where the conclusion of his studies happened and then this n will be our right endpoint of the domain being talked about... And

myininaya (myininaya):

other hints: \[\text{ part B} \\ \text{ the y-intercept is when what coordinate equals 0}? \\ \] \[\text{ part C } \text{ average rate of change from } x=a \text{ to } x=b \text { can be found by } \\ \text{ computing } \frac{f(b)-f(a)}{b-a}\]

OpenStudy (anonymous):

Thank you so much